# main_arXiv.bib

### v1

@@ Abstract
In this paper we analyze the sample complexities of learning the optimal state-action value function $Q^{*}$ and an optimal policy $\pi^{*}$ in a discounted Markov decision process (MDP) where the agent has recursive entropic risk-preferences with risk-parameter $\beta\neq 0$ and where a generative model of the MDP is available. We provide and analyze a simple model based approach which we call model-based risk-sensitive $Q$-value-iteration (MB-RS-QVI) which leads to $(\epsilon,\delta)$-PAC-bounds on $\|Q^{*}-Q^{k}\|$, and $\|V^{*}-V^{\pi_{k}}\|$ where $Q_{k}$ is the output of MB-RS-QVI after k iterations and $\pi_{k}$ is the greedy policy with respect to $Q_{k}$. Both PAC-bounds have exponential dependence on the effective horizon $\frac{1}{1-\gamma}$ and the strength of this dependence grows with the learners risk-sensitivity $|\beta|$. We also provide two lower bounds which shows that exponential dependence on $|\beta|\frac{1}{1-\gamma}$ is unavoidable in both cases. The lower bounds reveal that the PAC-bounds are both tight in $\varepsilon$ and $\delta$ and that the PAC-bound on $Q$-learning is tight in the number of actions $A$, and that the PAC-bound on policy-learning is nearly tight in $A$.
@@ 1 Introduction
In reinforcement learning (RL), the aim of the agent is to conventionally optimizes the expected return, which is defined in terms of a (discounted) sum of rewards [sutton1998reinforcement]. In the majority of RL literature, the environment is modeled via the Markov Decision Process (MDP) framework [puterman2014markov], wherein efficient computation of an optimal policy, thanks to optimal Bellman equations, renders possible. However, as a risk-neutral objective, the expected return fails to capture the true needs of many high-stake applications arising in, e.g., medical treatment [ernst2006clinical], finance [scutella2013robust, bielecki1999risk], and operations research [delage2010percentile]. Decision making in such applications must take into account the variability of returns, and risks thereof. To account for this, one may opt to to maximize a risk measure of the return distribution, while another approach could be to consider the entire distribution of return, as is done under the distributional RL framework [bellemare_distributional_2023] that has received much attention over the last decade.
@@
Within the first approach, the risk is quantified via concave risk measures, which yield well-defined mathematical optimization frameworks. Notably, they include value-at-risk (VaR), Conditional VaR (CVaR) [shapiro2021LNs], entropic risk [howard1972risk], and entropic VaR (EVaR) [ahmadi2012entropic], all of which have been applied to a wide-range of scenarios. CVaR appears to be the most popular one used to model risk-sensitivity in MDPs [chow2014algorithms, bisi2022risk, brown2020bayesian, bauerle_2011_markov], mainly due to a delicate control it offers for the undesirable tail of return. Despite its popularity and rich interpretation, solving and learning MDPs with CVaR-defined objectives has rendered technically challenging [bauerle_2011_markov]. This has been a key motivation of adopting weaker notions such as nested CVaR [bauerle_markov_2022], at the expense of sacrificing the interpretability. Entropic risk, as another popular notion, has been considered for risk control in MDPs [borkar2002risk, neu2017unified, hau2023entropic, hu2023tighter, fei2020risk]. In the RL literature, it has been mainly considered for the undiscounted settings, despite the popularity of discounted MDPs. A notable exception is [hau2023entropic] that studies, among other things, planning in discounted MDPs under the entropic risk.
@@
In this paper, we study risk-sensitive discounted RL where the agent’s objective is formulated using the entropic risk measure. In discounted RL, where future rewards are discounted by a factor of $\gamma$, one may identify two main approaches to apply the entropic risk to sequence of rewards, $(r_{t})_{t\geq 0}$, collected by an RL agent. The first and most intuitive one, which we may call the non-recursive approach, consists in directly applying the entropic risk functional to return $\sum_{t=0}^{\infty}\gamma^{t}r_{t}$. The other approach, which we may call the recursive approach, works by applying the risk functional at every step $t$ (see Section 2 for details). The non-recursive approach (e.g., [hau2023entropic]), while being most intuitive, has several drawbacks; e.g., the optimal policy might not be time-consistent (see [jaquette1976utility]). In contrast, the non-recursive approach yields a form of Bellman optimality equation, which is key to developing learning algorithms with provable sample complexity. Therefore, we restrict attention to the RL problem defined using recursive risk-preferences.
@@ 1.1 Main Contributions
We study risk-sensitive RL in finite discounted MDPs under the non-recursively applied entropic risk measure, assuming that the agent is given access to a generative model of the environment. The agent’s learning performance is assessed via sample complexity defined as the total number $T$ of samples needed to learn, for input $(\varepsilon,\delta)$, either an $\varepsilon$-optimal policy (which we call policy learning), or an $\varepsilon$-close approximation (in max-norm) to the optimal Q-value (which we call Q-value learning), with probability exceeding $1-\delta$.
Specifically, we make the following contributions: We present an algorithm called Model-Based Risk-Sensitive Q-Value Iteration (MB-RS-QVI), a model-based RL algorithms for the RL problem considered, which is derived using a simple plug-in estimator. It is provably sample efficient, despite its simple design. Notably, we report sample complexity bounds on the performance of (MB-RS-QVI) under Q-value learning (Theorem 1) and policy learning (Theorem 2), both scaling as $\widetilde{\mathcal{O}}\Big{(}\frac{S^{2}A}{\varepsilon^{2}(1-\gamma)^{2}|\beta|^{2}}e^{2|\beta|\frac{1}{1-\gamma}}\Big{)}$, in any discounted MDP with $S$ states, $A$ action, and discount factor $\gamma$, with $\widetilde{\mathcal{O}}$ hiding logarithmic factors. Here, $\beta$ denotes the risk parameter (see Section 2 for precise definition), where $\beta>0$ (resp. $\beta<0$) corresponds to a risk-averse (resp. risk-seeking) agent. An interesting property of these bounds is the exponential dependence on the effective horizon $\frac{1}{1-\gamma}$, which is absent in the conventional (risk-neutral) RL, wherein $\beta=0$. Another key contribution of this paper is to derive the first, to our knowledge, sample complexity lower bounds for the discounted RL under entropic risk measure. Our lower bounds establish that for both Q-value learning (Theorem 3) and policy learning (Theorem 4), one needs at least $\widetilde{\Omega}\Big{(}\frac{SA}{\varepsilon^{2}|\beta|^{2}}e^{|\beta|\frac{1}{1-\gamma}}\Big{)}$ samples to come $\varepsilon$-close to optimality. Interestingly, the derived lower bounds assert that exponential dependence on $\frac{|\beta|}{1-\gamma}$ in both sample complexities are unavoidable.
@@ 1.2 Related Work
There is a large body of papers studying provably-sample efficient learning algorithms in discounted MDPs. These papers consider a variety of settings, including the generative setting [kakade2003sample, gheshlaghi2013minimax, agarwal_model-based_2020], the offline (or batch) setting [rashidinejad2021bridging, li2024settling], and the online setting [strehl_analysis_2008, lattimore_near-optimal_2014]. In particular, in the generative setting – which we consider in this paper – some notable developments include, but not limited to [kearns1998phased_Q_learning, gheshlaghi2013minimax, agarwal_model-based_2020, li2020breaking, sidford2018variance, wang2020randomized, chen2020finite]. Among these, [gheshlaghi2013minimax, agarwal_model-based_2020, sidford2018variance, li2020breaking] present algorithms, attaining minimax-optimal sample complexity bounds, although some of these result do not cover the full range of $\varepsilon$. We also mention a line of work, comprising e.g. [panaganti2022sample, shi2023curious], that investigate discounted RL in the generative setting but under distributional robustness. Let us also remark that some recent works – notably [zurek2024span, zurek2025plugin] – study sample complexity of average-reward MDPs in the generative setting.
Finally, we mention that some studies consider adaptive sampling in the generative setting to account for the heterogeneity across the various state-action pairs in the MDP; see, e.g., [almarjani2021adaptive, zanette2019almost]. This line of works stand in contrast to the papers cited above that strive for optimizing the performance, in the worst-case sense, via uniformly sampling various state-action pairs.
@@
There exists a rich literature on decision making under a risk measure. We refer to [sani2012risk, maillard2013robust, khajonchotpanya2021revised] for some developments in bandits, where the performance is assessed via regret. Extensions to episodic MDPs were pursued in a recent line of work, including [fei2020risk, liang2024regret, hu2023tighter], which establish near-optimal guarantees on regret. The two lines of work (on bandits and episodic MDPs) constitute the majority of work on risk-sensitive RL, thoroughly studying a variety of risk measures including CVaR, entropic risk, and entropic value-at-risk [ahmadi2012entropic, su2025risk].
Among the various risk measures, CVaR has arguably received a great attention; see, for instance, [deng2025near, du2023provably, chen2024provably]. For instance, [du2023provably, chen2024provably] study episodic RL in the regret setting and present algorithms with sub-linear regret in tabular MDPs ([du2023provably]) and under function approximation ([chen2024provably]), while [deng2025near] investigates the sample complexity analysis in the generative setting (similar to ours), We also provide a pointer to some work which deal with a class of measures called coherent risk measures that include CVaR as a special case. In this category, we refer to, for instance, [tamar2015policy], which studied policy gradient algorithms, and to [lam2022risk] that investigates regret minimization in the episodic setting and with function approximation. In the case of infinite-horizon setting, the entropic risk measure is mostly considered in the undiscounted (i.e., the average-reward) setting; examples include [borkar2002risk, neu2017unified, marthe2023beyond, marthe2025efficient]. In contrast, little attention is paid to the discounted setting, which is mainly due to technical difficulties caused by discounting. The work [hau2023entropic] is among the few papers investigating solving discounted MDPs under the entropic risk.
@@
For $n\in\mathbb{N}$, let $[n]:=\{1,\ldots,n\}$. $1_{A}$ denotes the indicator function of an event $A$. Given a set $\mathcal{X}$, $\Delta(\mathcal{X})$ denotes the probability simplex over $\mathcal{X}$. We use the convention that $\|\cdot\|:=\|\cdot\|_{\infty}$ and explicitly use the subscript $\|\cdot\|_{p}$ when using $p$-norms for which $1\leq p<\infty$. We use $Z=\mathcal{S}\times\mathcal{A}$ to denote the set of all state-action pairs.
@@ 2.1 Markov Decision Processes
We write the 6-tuple $M=(\mathcal{S},\mathcal{A},P,R,\gamma,\beta)$ as a finite, discounted infinite-horizon Markov decision process (MDP), where $\mathcal{S}=\{1,2,...,S\}$ is the finite state space of size $S:=|\mathcal{S}|$, $\mathcal{A}=\{1,2,...,A\}$ is the finite action space of size $A:=|\mathcal{A}|$, $P:\mathcal{S}\times\mathcal{A}\rightarrow\Delta(\mathcal{S})$ is the transition probability function, $R:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ is the deterministic reward function, $\gamma\in(0,1)$ is the discount factor, and $\beta\neq 0$ is the risk-parameter. The agent interacts with the MDP $M$ as follows. At the beginning of the process, $M$ is in some initial state $s_{0}\in\mathcal{S}$. At each time $t\geq 0$, the agent is in state $s_{t}\in\mathcal{S}$ and decides on an action $a_{t}\in\mathcal{A}$ according to some rule. The MDP generates a reward $r_{t}:=R(s_{t},a_{t})$ and a next-state $s_{t+1}\sim P(\cdot|s_{t},a_{t})$. The MDP moves to $s_{t+1}$ when the next time slot begins, and this process continues ad infinitum. This process yields a growing sequence $(s_{t},a_{t},r_{t})_{t\geq 0}$. The agent’s goal is to maximize an objective function, as function of reward values $(r_{t})_{t\geq 0}$ in expectation, which involves the two parameters $\gamma$ and $\beta$; in addition to the discount factor $\gamma$ that makes future rewards less valuable than the present ones, the risk-parameter $\beta$ quantifies to what degree the agent seeks or avoids strategies that have more variability in the rewards obtained over time. To concretely define the objective function of the agent, we shall introduce some necessary concepts.
@@ 2.2 Entropic Risk Preferences
The entropic risk preferences can be justified by expected utility theory. Consider for $\beta\neq 0$ the class of utility functions $u(t)=\frac{1}{\beta}(1-e^{-\beta t})$ defined for $t\in\mathbb{R}$. The utility $u$ is supposed to describe the preferences of some economic agent in the form of how much utility $u(t)$ she derives from some monetary quantity $t\in\mathbb{R}$. For any bounded random variable $X\in L^{\infty}(\Omega,\mathcal{F},\mathbb{P})$, the certainty equivalent $u^{-1}(\mathbb{E}[u(X)])=\frac{-1}{\beta}\log(\mathbb{E}[e^{-\beta X}])$ expresses the amount of money that would give the same utility as that of entering in the bet given by the random variable $X$. We thus define the functional $\Phi_{\beta}:L^{\infty}(\Omega,\mathcal{F},\mathbb{P})\rightarrow\mathbb{R}$ by
$\Phi_{\beta}(X)=-\frac{1}{\beta}\log(\mathbb{E}[e^{-\beta X}])$
@@
We note that there does not seem to be consensus on wether the functional is parametrized with $\beta$ of $-\beta$. We follow this convention considering its widespread use in the actuarial literature [asienkiewicz2017note], but we remark that some other lines of work appear to prefer the other parametrization. For the case $\beta\rightarrow 0$, we recover the risk-neutral case that is simply the expectation: $\lim_{\beta\to 0}\Phi_{\beta}(X)=\mathbb{E}[X]$. It is straightforward to see that $\Phi_{\beta}$ admits the following properties:
$\Phi_{\beta}(X)\leq\Phi_{\beta}(Y),\quad\text{for any }X\leq Y,$ (1)
$\Phi_{\beta}(c)=c,\quad\text{for any }c\in\mathbb{R},$ (2)
$\Phi_{\beta}(X)\leq\mathbb{E}[X],\quad\text{for }\beta>0,$ (3)
$\Phi_{\beta}(X)\geq\mathbb{E}[X]\quad\text{for }\beta<0,$ (4)
@@
where properties (3)-(4) follow from Jensen’s inequality. Using $\Phi_{\beta}$ as a measure of the preference for different random variables, it follows directly from (2)-(4) that $\Phi_{\beta}(X)\leq\Phi_{\beta}(\mathbb{E}[X])$ for $\beta>0$ and $\Phi_{\beta}(X)\geq\Phi_{\beta}(\mathbb{E}[X])$ for $\beta<0$, which shows that $\beta>0$ is associated with risk-aversion, while $\beta<0$ is associated with risk-seeking behavior.
@@
Applying risk in a stochastic dynamic process can be done in several ways and is thus more complicated than for a single-period problem. Two approaches to this end exist in the literature. The first and most intuitive one, which we name the non-recursive approach, is to apply the functional to the total discounted sum of rewards $\Phi_{\beta}\big{(}\sum_{t=0}^{\infty}\gamma^{t}r_{t}\big{)}$, which is well-defined under the bounded rewards assumption, i.e., $r_{t}\in[0,1]$ for all $t$. The other approach, which we may call the recursive approach, works by applying the risk at every step $t$ where we give the details below. The non-recursive approach is probably most intuitive but has several drawbacks. Even though there is no obvious optimality equation for the non-recursive case, a solution to the planning problem has been proposed in [hau2023entropic]. In comparison, the planning problem is more straightforward with recursive risk-preferences due to the availability of an optimality equation that allows for simple value-iteration type algorithms. The recursive approach also guarantees the existence of an optimal stationary deterministic policy whereas with non-recursive risk preferences the optimal policy might not be time-consistent (see [jaquette1976utility]). In this paper, we study the problem with recursive risk-preferences.

### v2

# Entropic Risk Optimization in Discounted MDPs: Sample Complexity Bounds with a Generative Model

@@ Abstract
In this paper, we analyze the sample complexities of learning the optimal state-action value function $Q^{*}$ and an optimal policy $\pi^{*}$ in a finite discounted Markov decision process (MDP) where the agent has recursive entropic risk-preferences with risk-parameter $\beta\neq 0$ and where a generative model of the MDP is available. We provide and analyze a simple model based approach which we call model-based risk-sensitive $Q$-value-iteration (MB-RS-QVI) which leads to $(\varepsilon,\delta)$-PAC-bounds on $\|Q^{*}-Q^{k}\|$, and $\|V^{*}-V^{\pi_{k}}\|$ where $Q_{k}$ is the output of MB-RS-QVI after k iterations and $\pi_{k}$ is the greedy policy with respect to $Q_{k}$. Both PAC-bounds have exponential dependence on the effective horizon $\frac{1}{1-\gamma}$ and the strength of this dependence grows with the learners risk-sensitivity $|\beta|$. We also provide two lower bounds which shows that exponential dependence on $|\beta|\frac{1}{1-\gamma}$ is unavoidable in both cases. The lower bounds reveal that the PAC-bounds are tight in the parameters $S,A,\delta,\varepsilon$ and that unlike in the classical setting it is not possible to have polynomial dependence in all model parameters.
@@ 1 Introduction
In reinforcement learning (RL), the aim of the agent is to conventionally optimize the expected return, which is defined in terms of a (discounted) sum of rewards sutton1998reinforcement . In the majority of RL literature, the environment is modeled via the Markov Decision Process (MDP) framework puterman2014markov , wherein efficient computation of an optimal policy, thanks to optimal Bellman equations, renders possible. However, as a risk-neutral objective, the expected return fails to capture the true needs of many high-stake applications arising in, e.g., medical treatment ernst2006clinical , finance scutella2013robust ; bielecki1999risk , and operations research delage2010percentile . Decision making in such applications must take into account the variability of returns, and risks thereof. To account for this, one may opt to to maximize a risk measure of the return distribution, while another approach could be to consider the entire distribution of return, as is done under the distributional RL framework bellemare_distributional_2023 that has received much attention over the last decade.
@@
Within the first approach, the risk is quantified via concave risk measures, which yield well-defined mathematical optimization frameworks. Notably, they include value-at-risk (VaR), Conditional VaR (CVaR) shapiro2021LNs , entropic risk howard1972risk , and entropic VaR (EVaR) ahmadi2012entropic , all of which have been applied to a wide-range of scenarios. CVaR appears to be the most popular one used to model risk-sensitivity in MDPs chow2014algorithms ; bisi2022risk ; brown2020bayesian ; bauerle_2011_markov , mainly due to a delicate control it offers for the undesirable tail of return. Despite its popularity and rich interpretation, solving and learning MDPs with CVaR-defined objectives has rendered technically challenging bauerle_2011_markov . This has been a key motivation of adopting weaker notions such as nested CVaR bauerle_markov_2022 , at the expense of sacrificing the interpretability. Entropic risk, as another popular notion, has been considered for risk control in MDPs borkar2002risk ; neu2017unified ; hau2023entropic ; hu2023tighter ; fei2020risk . In the RL literature, it has been mainly considered for the undiscounted settings, despite the popularity of discounted MDPs. A notable exception is hau2023entropic that studies, among other things, planning in discounted MDPs under the entropic risk.
@@
In this paper, we study risk-sensitive discounted RL where the agent’s objective is formulated using the entropic risk measure. In discounted RL, where future rewards are discounted by a factor of $\gamma$, one may identify two main approaches to apply the entropic risk to sequence of rewards, $(r_{t})_{t\geq 0}$, collected by an RL agent. The first and most intuitive one, which we may call the non-recursive approach, consists in directly applying the entropic risk functional to return $\sum_{t=0}^{\infty}\gamma^{t}r_{t}$. The other approach, which we may call the recursive approach, works by applying the risk functional at every step $t$ (see Section 2 for details). The non-recursive approach (e.g., hau2023entropic ), while being most intuitive, has several drawbacks; e.g., the optimal policy might not be time-consistent (see jaquette1976utility ). In contrast, the recursive approach yields a form of Bellman optimality equation, which is key to developing learning algorithms with provable sample complexity. Therefore, we restrict attention to the RL problem defined using recursive risk-preferences.
@@ 1.1 Main Contributions
We study risk-sensitive RL in finite discounted MDPs under the recursively applied entropic risk measure, assuming that the agent is given access to a generative model of the environment. The agent’s learning performance is assessed via sample complexity defined as the total number $T$ of samples needed to learn, for input $(\varepsilon,\delta)$, either an $\varepsilon$-optimal policy (which we call policy learning), or an $\varepsilon$-close approximation (in max-norm) to the optimal Q-value (which we call Q-value learning), with probability exceeding $1-\delta$.
Specifically, we make the following contributions: We present an algorithm called Model-Based Risk-Sensitive Q-Value Iteration (MB-RS-QVI), a model-based RL algorithms for the RL problem considered, which is derived using a simple plug-in estimator. It is provably sample efficient, despite its simple design. Notably, we report sample complexity bounds on the performance of (MB-RS-QVI) under Q-value learning (Theorem 1) scaling as $\widetilde{O}\Big(\frac{SA}{\varepsilon^{2}(1-\gamma)^{2}}(\frac{e^{|\beta|\frac{1}{1-\gamma}}-1}{|\beta|})^{2}\Big)$ and policy learning (Theorems 2 and 3), scaling as $\widetilde{O}\Big(\frac{SA}{\varepsilon^{2}(1-\gamma)^{4}}(\frac{e^{|\beta|\frac{1}{1-\gamma}}-1}{|\beta|})^{2}\Big)$, and $\widetilde{O}\Big(\frac{S^{2}A}{\varepsilon^{2}(1-\gamma)^{2}}\frac{e^{|\beta|\frac{1}{1-\gamma}}}{|\beta|^{2}}\Big)$ in any discounted MDP with $S$ states, $A$ action, and discount factor $\gamma$, with $\widetilde{\mathcal{O}}$ hiding logarithmic factors. Here, $\beta$ denotes the risk parameter (see Section 2 for precise definition), where $\beta>0$ (resp. $\beta<0$) corresponds to a risk-averse (resp. risk-seeking) agent. An interesting property of these bounds is the exponential dependence on the effective horizon $\frac{1}{1-\gamma}$, which is absent in the conventional (risk-neutral) RL, wherein $\beta=0$. Another key contribution of this paper is to derive the first, to our knowledge, sample complexity lower bounds for the discounted RL under entropic risk measure. Our lower bounds establish that for Q-value learning (Theorem 4) and policy learning (Theorem 5), one needs at least $\widetilde{\Omega}\Big(\frac{(S-2)A}{\varepsilon^{2}}\frac{e^{|\beta|\frac{1}{1-\gamma}}-3}{|\beta|^{2}}\Big)$ and $\widetilde{\Omega}\Big(\frac{(S-2)(A-1)}{\varepsilon^{2}}\frac{e^{|\beta|\frac{1}{1-\gamma}}-3}{|\beta|^{2}}\Big)$ samples, respectively, to come $\varepsilon$-close to optimality. Interestingly, the derived lower bounds assert that exponential dependence on $\frac{|\beta|}{1-\gamma}$ in both sample complexities are unavoidable thus making the risk-sensitive setting fundamentally harder than the classical risk-neutral setting.
@@ 1.2 Related Work
There is a large body of papers studying provably-sample efficient learning algorithms in discounted MDPs. These papers consider a variety of settings, including the generative setting kakade2003sample ; gheshlaghi2013minimax ; agarwal_model-based_2020 , the offline (or batch) setting rashidinejad2021bridging ; li2024settling , and the online setting strehl_analysis_2008 ; lattimore_near-optimal_2014 . In particular, in the generative setting – which we consider in this paper – some notable developments include, but not limited to kearns1998phased_Q_learning ; gheshlaghi2013minimax ; agarwal_model-based_2020 ; li2020breaking ; sidford2018variance ; wang2020randomized ; chen2020finite . Among these, gheshlaghi2013minimax ; agarwal_model-based_2020 ; sidford2018variance ; li2020breaking present algorithms attaining minimax-optimal sample complexity bounds, although some of these result do not cover the full range of $\varepsilon$. We also mention a line of work, comprising e.g. panaganti2022sample ; shi2023curious , that investigate discounted RL in the generative setting but under distributional robustness. Let us also remark that some recent works – notably zurek2024span ; zurek2025plugin – study sample complexity of average-reward MDPs in the generative setting.
Finally, we mention that some studies consider adaptive sampling in the generative setting to account for the heterogeneity across the various state-action pairs in the MDP; see, e.g., almarjani2021adaptive ; zanette2019almost . This line of works stand in contrast to the papers cited above that strive for optimizing the performance, in the worst-case sense, via uniformly sampling various state-action pairs.
@@
There exists a rich literature on decision making under a risk measure. We refer to sani2012risk ; maillard2013robust ; khajonchotpanya2021revised for some developments in bandits, where the performance is assessed via regret. Extensions to episodic MDPs were pursued in a recent line of work, including fei2020risk ; liang2024regret ; hu2023tighter , which establish near-optimal guarantees on regret. The two lines of work (on bandits and episodic MDPs) constitute the majority of work on risk-sensitive RL, thoroughly studying a variety of risk measures including CVaR, entropic risk, and entropic value-at-risk ahmadi2012entropic ; su2025risk .
Among the various risk measures, CVaR has arguably received a great attention; see, for instance, deng2025near ; du2023provably ; chen2024provably . For instance, du2023provably ; chen2024provably study episodic RL in the regret setting and present algorithms with sub-linear regret in tabular MDPs (du2023provably ) and under function approximation (chen2024provably ), while deng2025near investigates the sample complexity analysis in the generative setting (similar to ours), We also provide a pointer to some work which deal with a class of measures called coherent risk measures that include CVaR as a special case. In this category, we refer to, for instance, tamar2015policy , which studies policy gradient algorithms, and to lam2022risk that investigates regret minimization in the episodic setting and with function approximation. In the case of infinite-horizon setting, the entropic risk measure is mostly considered in the undiscounted (i.e., the average-reward) setting; examples include borkar2002risk ; neu2017unified ; marthe2023beyond ; marthe2025efficient . In contrast, little attention is paid to the discounted setting, which is mainly due to technical difficulties caused by discounting. The work hau2023entropic is among the few papers investigating solving discounted MDPs under the entropic risk.
@@
For $n\in\mathbb{N}$, let $n]:=\{1,\ldots,n\}$. $1_{A}$ denotes the indicator function of an event $A$. Given a set $\mathcal{X}$, $\Delta(\mathcal{X})$ denotes the probability simplex over $\mathcal{X}$. We use the convention that $\|\cdot\|:=\|\cdot\|_{\infty}$ and explicitly use the subscript $\|\cdot\|_{p}$ when using $p$-norms for which $1\leq p<\infty$. We use $Z=\mathcal{S}\times\mathcal{A}$ to denote the set of all state-action pairs.
@@ 2.1 Markov Decision Processes
We write the 6-tuple $M=(\mathcal{S},\mathcal{A},P,R,\gamma,\beta)$ as a finite, discounted infinite-horizon Markov decision process (MDP), where $\mathcal{S}=\{1,2,...,S\}$ is the finite state space of size $S:=|\mathcal{S}|$, $\mathcal{A}=\{1,2,...,A\}$ is the finite action space of size $A:=|\mathcal{A}|$, $P:\mathcal{S}\times\mathcal{A}\rightarrow\Delta(\mathcal{S})$ is the transition probability function, $R:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ is the deterministic reward function, $\gamma\in(0,1)$ is the discount factor, and $\beta\neq 0$ is the risk-parameter. The agent interacts with the MDP $M$ as follows. At the beginning of the process, $M$ is in some initial state $s_{0}\in\mathcal{S}$. At each time $t\geq 0$, the agent is in state $s_{t}\in\mathcal{S}$ and decides on an action $a_{t}\in\mathcal{A}$ according to some rule. The MDP generates a reward $r_{t}:=R(s_{t},a_{t})$ and a next-state $s_{t+1}\sim P(\cdot|s_{t},a_{t})$. The MDP moves to $s_{t+1}$ when the next time slot begins, and this process continues ad infinitum. This process yields a growing sequence $(s_{t},a_{t},r_{t})_{t\geq 0}$. The agent’s goal is to maximize an objective function, as a function of the reward values $(r_{t})_{t\geq 0}$ which involves the two parameters $\gamma$ and $\beta$; in addition to the discount factor $\gamma$ that makes future rewards less valuable than the present ones, the risk-parameter $\beta$ quantifies to what degree the agent seeks or avoids strategies that have more variability in the rewards obtained over time. To concretely define the objective function of the agent, we shall introduce some necessary concepts.
@@ 2.2 Entropic Risk Preferences
The entropic risk preferences is rooted in expected utility theory. Consider for $\beta\neq 0$ the class of utility functions $u(t)=\frac{1}{\beta}(1-e^{-\beta t})$ defined for $t\in\mathbb{R}$. The utility $u$ is supposed to describe the preferences of some economic agent in the form of how much utility $u(t)$ she derives from some monetary quantity $t\in\mathbb{R}$. For any bounded random variable $X\in L^{\infty}(\Omega,\mathcal{F},\mathbb{P})$, the certainty equivalent $u^{-1}(\mathbb{E}[u(X)])=\frac{-1}{\beta}\log(\mathbb{E}[e^{-\beta X}])$ expresses the amount of money that would give the same utility as that of entering in the bet given by the random variable $X$. We thus define the functional $\rho:L^{\infty}(\Omega,\mathcal{F},\mathbb{P})\rightarrow\mathbb{R}$ by
$\rho(X)=-\frac{1}{\beta}\log(\mathbb{E}[e^{-\beta X}])\,.$
@@
We note that there does not seem to be consensus on wether the functional is parametrized with $\beta$ of $\beta$. We follow this convention considering its widespread use in the actuarial literature asienkiewicz2017note , but we remark that some other lines of work appear to prefer the other parametrization. For the case $\beta\rightarrow 0$, we recover the risk-neutral case that is simply the expectation: $\lim_{\beta\to 0}\rho(X)=\mathbb{E}[X]$. It is straightforward to see that $\rho$ admits the following properties:
$\rho(X)\leq\rho(Y),\quad\text{for any }X\leq Y,$ (1)
$\rho(c)=c,\quad\text{for any }c\in\mathbb{R},$ (2)
$\rho(X)\leq\mathbb{E}[X],\quad\text{for }\beta>0,$ (3)
$\rho(X)\geq\mathbb{E}[X]\quad\text{for }\beta<0,$ (4)
@@
where properties (3)-(4) follow from Jensen’s inequality. Using $\rho$ as a measure of the preference for different random variables, it follows directly from (2)-(4) that $\rho(X)\leq\rho(\mathbb{E}[X])$ for $\beta>0$ and that $\rho(X)\geq\rho(\mathbb{E}[X])$ for $\beta<0$. If further shows that $\beta>0$ is associated with risk-aversion, while $\beta<0$ is associated with risk-seeking behavior.
@@
Applying risk in a stochastic dynamic process can be done in several ways and is thus more complicated than for a single-period problem. Two approaches to this end exist in the literature. The first and most intuitive one, which is often called the static or non-recursive approach, is to apply the functional to the total discounted sum of rewards $\rho\big(\sum_{t=0}^{\infty}\gamma^{t}r_{t}\big)$, which is well-defined under the bounded rewards assumption, i.e., $r_{t}\in[0,1]$ for all $t$. The other approach, which we may call the recursive approach, works by applying the risk at every step $t$ where we give the details below. The non-recursive approach is probably most intuitive but has several drawbacks. Even though there is no obvious optimality equation for the non-recursive case, a solution to the planning problem has been proposed in hau2023entropic . In comparison, the planning problem is more straightforward with recursive risk-preferences due to the availability of an optimality equation that allows for simple value-iteration type algorithms. The recursive approach also guarantees the existence of an optimal stationary deterministic policy whereas with non-recursive risk preferences the optimal policy might not be time-consistent (see jaquette1976utility ). In this paper, we study the problem with recursive risk-preferences.